# Generalized Chern-Pontryagin models

### v2

@@ IV Generating the Chern-Simons term
We start this section by showing that generalized Chern-Pontryagin models admit both Jordan and Einstein frame representations similar to what happens with $f(R)$ theories DeFelice . To see that more clearly, it is useful to observe that the action (1) can be cast into the following form
@@
is the interacting potential between both scalar fields. Note that the first and last terms in the gravity sector of the action (42) resemble a BD action with the difference that the potential is composed of the BD scalar $\Phi$ field, in addition to the scalar field $\vartheta$. The second term in Eq. (42) is just the topological Chern-Pontryagin term non-minimally coupled with $\vartheta$ (this term is frequently called the Chern-Simons one, see f.e. JaPi ). Therefore, starting from the action (1), we find a natural manner, different from the original proposal JaPi , to generate the Chern-Simons term, with $\vartheta$ playing the role of the CS coupling field Yunes and being one of the two scalar degrees of freedom coming up in the scalar-tensor representation. In this sense, the action of the model written in the Jordan frame representation can be thought of as a generalization of the Chern-Simons modified gravity. Note, however, that the previous choice, $f(R,\,^{*}RR)=R+\beta(\,^{*}RR)^{2}$, does not admit a scalar-tensor representation since the on-shell condition (39) is not ensured because (40) does not hold.
@@
where the tilde describes geometrical quantities in the Einstein frame and $\Omega$ is the conformal factor. The Ricci scalar in both frames is related to each other by
@@
$\tilde{S}$ $=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\tilde{g}}\left\{\tilde{R}+\vartheta\,^{*}\tilde{R}\tilde{R}-\frac{1}{4e^{2\sqrt{\frac{2}{3}}\kappa\psi}}\left[\frac{\vartheta^{2}}{\beta}+\frac{1}{\alpha}\left(e^{\sqrt{\frac{2}{3}}\kappa\psi}-1\right)^{2}\right]\right\}+$ (51)
$+$ $S_{m}(e^{-\sqrt{\frac{2}{3}}\kappa\psi}\tilde{g}_{\mu\nu},\Psi),$
which is the non-dynamical Chern-Simons modified gravity plus a non-trivial potential depending on two interacting scalar fields.
IV.1 The field equations in the Einstein frame
@@
IV.1.1 Gödel-type vacuum solutions in the presence of a trivial potential $U=0$.
In this case, the field equations (added by a cosmological constant) reduce to a similar shape to those found in Godelour2 unless by the fact that, here, $\tilde{g}$ is the dynamical metric. By solving the field equations, the authors of Godelour2 found the following solutions for the scalar fields: $\vartheta=b(z-z_{0})$ and $\psi=s(z-z_{0})$, where $z_{0}$ is an arbitrary real constant. Therefore, formally, the case corresponding to the trivial potential, $U=0$, just reduces to those found for CSMG Godelour1 ; Godelour2 ; Godelour3 , as expected.
@@ V Summary
It is easy to check that many known metrics solve the equations of motion for our model, namely, all those ones which yield the result ${}^{*}RR=0$; in this case, the field equations reduce to Einstein ones. In particular, it is true for all spherically symmetric metrics. This fact motivated us to study perturbations of two well-known metrics, namely, Schwarzschild and Gödel ones. In a certain sense, our study can be treated as a continuation of the paper Alt where the perturbations of the Gödel-type metric have been considered within the non-dynamical CSMG. We investigated the consistency of these metrics within the particular model $f(R,{}^{*}RR)=R+\beta({}^{*}RR)^{2}$. For the perturbations of the Schwarzschild black hole, we have found a first-order frame-dragging effect in the perturbation parameter $L=\beta M^{2}$, apart from the standard GR term. However, for the perturbations of the Gödel-type metrics, we concluded that all first-order perturbations are equal to zero to achieve consistency of the field equations.
Similar to $f(R)$ theories, generalized Chern-Pontryagin models admit a two-field scalar-tensor representation with an interacting potential of such two scalar fields and, as a consequence, Jordan and Einstein frames. In this representation, it is clear the link between these models and Chern-Simons modified gravity. Actually, by describing the theory in the Einstein frame, we have shown that the action of these models is to some extent an intermediate case between the non-dynamical and dynamical CSMG. Furthermore, we have succeeded in engendering the Chern-Simons term more naturally than in the original proposal of CSMG JaPi .
The natural continuation of this study could consist in the study of cosmological perturbations and also the implications of the parity violation in the context of gravitational waves. We plan to address these issues in the forthcoming paper.

### v3

@@ IV Generating the Chern-Simons term
We start this section by showing that generalized Chern-Pontryagin models admit both Jordan and Einstein frame representations similar to what happens with $f(R)$ theories DeFelice .
IV.1 Jordan frame representation
To see how the Jordan frame representation emerges in more details, it is useful to observe that the action (1) can be cast into the following form
@@ IV.1 Jordan frame representation
is the interacting potential between both scalar fields. Note that the first and last terms in the gravity sector of the action (45) resemble a BD action with the difference that the potential is composed of the BD scalar $\Phi$ field (scalaron), in addition to the scalar field $\vartheta$. The second term in Eq. (45) is just the topological Chern-Pontryagin term non-minimally coupled with $\vartheta$ (this term is frequently called the Chern-Simons one, see f.e. JaPi ). Therefore, starting from the action (1), we find a natural manner, different from the original proposal JaPi , to generate the Chern-Simons term, with $\vartheta$ playing the role of the CS coupling field Yunes and being one of the two scalar degrees of freedom coming up in the scalar-tensor representation. In this sense, the action of the model written in the Jordan frame representation can be thought of as a generalization of the Chern-Simons modified gravity. Note, however, that the previous choice, $f(R,\,^{*}RR)=R+\beta(\,^{*}RR)^{2}$, does not admit a two-field scalar-tensor representation since the on-shell condition (42) is not ensured because (43) does not hold.
IV.1.1 The field equations in the Jordan frame representation
By varying the action (45) with respect to the metric and scalar fields: $g_{\mu\nu}$, $\Phi$ and $\vartheta$, we are able to find the following field equations in the Jordan frame representation, namely,
$G_{\mu\nu}$ $=$ $\frac{\kappa^{2}}{\Phi}(T^{(m)}_{\mu\nu}-g_{\mu\nu}V)+\frac{1}{\Phi}\left(\nabla_{\mu}\partial_{\nu}\Phi-g_{\mu\nu}\Box\Phi\right)-\frac{4}{\Phi}C_{\mu\nu};$ (47)
$R$ $=$ $2\kappa^{2}\frac{\partial V}{\partial\Phi};$ (48)
$\,{}^{*}RR$ $=$ $2\kappa^{2}\frac{\partial V}{\partial\vartheta},$ (49)
where we defined the Cotton tensor as
$C_{\mu\nu}=v_{\alpha}\epsilon^{\alpha}_{\,\,\,\beta\sigma_{(}\mu}\nabla^{\sigma}R^{\beta}_{\,\,\,\nu)}+v_{\alpha\beta}\,^{*}R^{\alpha\,\,\,\,\,\,\,\,\,\beta}_{\,\,\,(\mu\nu)},$ (50)
with $v_{\alpha}=\nabla_{\alpha}\vartheta$ and $v_{\alpha\beta}=\nabla_{\alpha}v_{\beta}$. Tracing Eq.(47), one finds
$\Box\Phi=\frac{1}{3}\Phi R+\frac{1}{3}\kappa^{2}(T^{(m)}-4V).$ (51)
Now, combining it with Eq.(48), we get
$\Box\Phi=\frac{1}{3}\Phi\frac{\partial V}{\partial\Phi}+\frac{1}{3}\kappa^{2}(T^{(m)}-4V),$ (52)
which means that the scalaron fulfills a dynamical field equation that depends on the trace of the stress-energy tensor and also the interacting potential between both scalar fields. On the other hand, the CS field $\vartheta$ satisfies a non-dynamical equation (49), whose solution provides an equation $\vartheta=\vartheta(\Phi)$, and then $\vartheta$ can be completely factored out (on-shell) from Eq.(45). Therefore, there exists only one scalar propagating degree of freedom in the theory.
We now take the model $f(R,\,^{*}RR)=R+\beta(\,^{*}RR)^{2}$ (the same one considered in the last section). In this case, the correspondent potential is simply given by
$V(\vartheta)=\frac{\vartheta^{2}}{8\beta\kappa^{2}},$ (53)
while the scalaron reduces to $\Phi=1$. By plugging this potential into Eq.(49), one finds
$\vartheta=2\beta(\,^{*}RR).$ (54)
which can be integrated out of Eq.(47) and, as a result, obtaining the field equations (2) for the aforementioned particular model. Therefore, as expected, one concludes that the field equations of the generalized Chern-Pontryagin gravity in the Jordan frame are equivalent to those in the standard representation.
IV.2 Einstein frame representation
@@
where the tilde describes geometrical quantities in the Einstein frame and $\Omega$ is the conformal factor (we note that this transformation is consistent only if $\partial f/\partial R>0$). The Ricci scalar in both frames is related to each other by
@@
$\tilde{S}$ $=$ $\frac{1}{2\kappa^{2}}\int d^{4}x\sqrt{-\tilde{g}}\left\{\tilde{R}+\vartheta\,^{*}\tilde{R}\tilde{R}-\frac{1}{4e^{2\sqrt{\frac{2}{3}}\kappa\psi}}\left[\frac{\vartheta^{2}}{\beta}+\frac{1}{\alpha}\left(e^{\sqrt{\frac{2}{3}}\kappa\psi}-1\right)^{2}\right]\right\}-$ (62)
-- $\frac{1}{2}\tilde{g}^{\mu\nu}\tilde{\nabla}_{\mu}\psi\tilde{\nabla}_{\nu}\psi+S_{m}(e^{-\sqrt{\frac{2}{3}}\kappa\psi}\tilde{g}_{\mu\nu},\Psi),$
which is the non-dynamical Chern-Simons modified gravity plus a non-trivial Lagrangian depending on two interacting scalar fields.
IV.2.1 The field equations in the Einstein frame
@@
IV.2.2 Gödel-type vacuum solutions in the presence of a trivial potential $U=0$
In the case of a trivial potential, the field equations (added by a cosmological constant) reduce to a similar shape to those found in Godelour2 unless by the fact that, here, $\tilde{g}$ is the dynamical metric. By solving the field equations, the authors of Godelour2 found the following solutions for the scalar fields: $\vartheta=b(z-z_{0})$ and $\psi=s(z-z_{0})$, where $z_{0}$ is an arbitrary real constant. Therefore, formally, the case corresponding to the trivial potential, $U=0$, just reduces to those found for CSMG Godelour1 ; Godelour2 ; Godelour3 , as expected.
@@ V Summary
It is easy to check that many known metrics solve the equations of motion for our model, namely, all those ones which yield the result ${}^{*}RR=0$; in this case, the field equations reduce to Einstein ones. In particular, it is true for all spherically symmetric metrics. This fact motivated us to study perturbations of two well-known metrics, namely, Schwarzschild and Gödel ones. In a certain sense, our study can be treated as a continuation of the paper Alt where the perturbations of Gödel-type metrics have been considered within the non-dynamical CSMG. We investigated the consistency of these metrics within the particular models $f(R,{}^{*}RR)=R+\beta({}^{*}RR)^{2}$ and $f(R,{}^{*}RR)=R+\alpha R^{2}+\beta({}^{*}RR)^{2}$. For the perturbations of the Schwarzschild black hole, we have found a first-order frame-dragging effect in the perturbation parameter $L=\beta M^{2}$, apart from the standard GR term. For the perturbations of the Gödel-type metrics, we have investigated both general first-order perturbations and those displaying axial symmetry breaking along the $z-$direction. We conclude that, in these scenarios, all perturbed metric functions must equal zero to ensure consistency with the field equations.
Similar to $f(R)$ theories, generalized Chern-Pontryagin models admit a two-field scalar-tensor representation with an interacting potential of such two scalar fields and, as a consequence, Jordan and Einstein frames. In this representation, it is clear the link between these models and Chern-Simons modified gravity. Actually, by describing the theory in the Einstein frame, we have shown that the action of this class of models is to some extent an intermediate case between the non-dynamical and dynamical CSMG. Furthermore, we have succeeded in engendering the Chern-Simons term more naturally than in the original proposal of CSMG JaPi .
The natural continuation of this study could involve investigating the study of cosmological perturbations and also the implications of the parity violation in the context of gravitational waves. We plan to address these issues in the forthcoming paper.